Can the fraction 54 represent a probability?
Since 5/4 > 1, it cannot represent a probability of an event A, denoted by Pr(A), because for any event A, 0 ≤ Pr(A) ≤ 1. If A never occur, then Pr(A) = 0; if A always occur, then Pr(A) = 1.
1/12 =0.08333... ≈ 8.3%
What is the probability that a human can walk on Jupiter?
---- It's absoloutly impossible. 3 reasons are that firstly the planet has an atmosphere that is made up from gases that are deadly poisonous to a human. Secondly, the surface of Jupiter (and just about all of the rest of it too) is gaseous and the only "solid" part is a very small inner core. Thirdly, even if you could find a solid surface to stand on you would still die instantly because the pressure on Jupiter is so incredibly heavy that a human would be crushed and flattened like a ribbon instantly.
If you toss three coins what is the probability that two coins will be heads up?
Possibilities: HHH, HHT, HTH, THH, HTT, THT, TTH, TTT. There are 3 chances out of 8 that there will be two heads and one more that there will be AT LEAST two heads.
What is the probability of drawing an ace from a deck of cards?
1 in 13
4 aces 4 divided by 4= 1
52 cards 52 divided by 4=13
What is the probability of a human offspring being female?
50% because the male and female would both have the chance
When you flip a coin what is the probability that it will come up tails?
the probability would be 50 to 50 chances
There's generally a 50% chance it will come up tails, but some coins have heavier designs on one side, so these may be more biased to a head or a tail over the term.
If it is a fair coin, then 0.5
1. to shoot, as with a rifle, at targets selected at whim: to plink at coins tossed in the air.
http://dictionary.reference.com/browse/plink
If there are 270725 possible 4-card hands?
The number of possible 4-card hands out of a 52 card deck is 270,725.
If an event is unlikely to happen its probability is?
If the even is never going to take place, and it is for sure, then it's probability will be 0.
For example, if the event is : A child being born with powers like Superman. For this event, the probability would be 0!
But in practicality, if the event is unlikely to happen, it's probability is likely to be nearer to 0. Keeping in mind the difference between "unlikely to happen" and "never going to happen for sure".
Because there is a possibility that the event that is unlikely to happen may just happen once in a thousand years. So in this case, the probability will be very very close to 0, but it will not be 0 exactly.
What is probability how does probability relate to genetics?
Probability is what chance something has to happen. The Punnett Square is a way how to predict in genetics how likely an offspring is to have a trait passed on from parents, or in other words find out the probability of a trait being in the phenotype or the genotype.
If the 1, 2, 3, 4 and 5 each appear in the numbers (no repeats, no eliminations-- so numbers like 11133 are not valid) then there are 120 combinations. If you can use any of the 5 any number of times up to a max of 5 times, (any possible way of making a 5 digit number using only 1, 2, 3, 4 and 5, so that 33455 for example is valid) then the number of combinations is 3125. The 3125 combinations include the 120 combinations above. This is a permutation of 5 objects, that can be thought of as an arrangement or a rearrangement of the five numbers. The number of permutations of this 5 numbers is equal to 5!. 5! = 5 x 4 x 3 x 2 x 1 = 120 ways. Or you can use the graphing calculator to compute 5!, such as: Press 5, MATH, with the right arrow go to PRB, press 4 (for !), ENTER.
13/36
You have a 1/9 chance of landing a 2 on the first spin and a 1/9 chance of landing 5 on the second, so the chances of landing on a 2 then a 5 should be (1/9)*(1/9) = 1/81
What does it mean to call a spade a spade?
It means to speak plainly.
Eg. : Would you like a piece of gum? - or - Would you like a mint?
- vs -
Your breath stinks.
What is a triangle with three sides and three angles?
Any shape with three sides and three angles is called a triangle.
odds in favor is when the number you want is first and the number out of is last.